OpenStudy (anonymous):

I need to find the slope, and x and y intercept of 3X-Y=-X+8

OpenStudy (anonymous):

-y=-4x+8 y=4x-8 slope is 4/1 and the y intercept is when x=0 so thats -8 the x intercept is when y=0 0=4x-8 8=4x x=2 (2,0) x intercept and (0,-8) y intercept

OpenStudy (anonymous):

thank you